# Robotic Sorting Systems: Robot Management and Layout Design Optimization

### v3

@@ 4.2 Concept of RC-S
The entry cycle of a specific feasible path is the earliest cycle at which all VPs in the path are unoccupied.
Algorithm 1 A heuristic method to implement RC-S
0:  Load the reservation table that records the status of VPs on the map within $N_{p}$ cycles. Set the maximum search range in each cycle as $N_{t}$. Calculate feasible path set $\mathcal{R}$.
@@
0:  At the beginning of cycle $t$, add the first robot of each queue at entrance to the candidate list.
0:  If the candidate list is empty, go to step 5; else, find the robot with the longest waiting time in the candidate list, obtain its origin $i$ and target outlet $k$. Calculate the entry cycle of each feasible path in $\mathcal{R}_{i,k}$, and choose the feasible path with the smallest entry cycle.
0:  If the entry cycle is less than $N_{t}$, reserve the VPs according to the feasible path, and assign the path to the robot; else, add the waiting time of the robot by $\tau_{c}$.
0:  Remove the robot from the candidate list and go to step 2.
@@
0:  Record the newly generated VPs in the reservation table, labeled as unoccupied. Remove the VPs which have left the network. Output the assignment of feasible path.
So far, we have identified the essential components required to integrate heuristics into the RC-S framework: recording the status of VPs ever existed, and for each delivery task, finding a feasible path with an entry cycle, and reserving the VPs in the space-time map. To improve robot transporting efficiency and ensure the accessibility of each outlet, we allow up to three turns in each delivery route under this scheme. Compared with classic MAPF methods, it avoids potential path conflicts from the design stage and greatly reduces the search range while maximizing intersection capacity. The optimal strategy of RC-S requires selecting a combination of feasible path and entry time for each robot, and checking the occupancy status of VPs.
Algorithm 1 has a complexity of $O(N_{t}\cdot\lvert R\rvert)$ in each cycle. Moreover, when the algorithm is forcibly terminated midway, it can still produce a feasible solution, where robots that are not served will wait until the next cycle. The performance of Algorithm 1 will be validated in section 6.
@@ 5 Performance Analysis of RC-S
Two important metrics in traffic analysis (Daganzo,, 2010) are the total vehicular distance traveled per second of operation $\bar{m}(veh\cdot m/s)$, and the average travel distance $\bar{l}$. The relationship among system throughput $\tilde{T}_{O}$ and these two metrics is shown in equation (11):
@@
Under the mechanism of RC-S, the first metric $\bar{m}$ is numerically consistent with the product of the number of occupied VPs, denoted by $n^{occupied}_{VP}$, and the speed of each VP, denoted by $v_{VP}$:
$\bar{m}=n^{occupied}_{VP}\cdot v_{VP}$ (12)
The speed of each VP could be obtained by expression (4)-(5) with the parameters of RC-S. For the number of occupied VPs, an intuitive idea is that it is related to the number of workers engaged in loading parcels in the loading zone and the spatial utilization under RC-S. We introduce two factors: workforce factor $\kappa$, representing the impact of the number of workers on the quantity of VPs that can be occupied; attenuation factor $\beta$, representing the utilization of VPs by the feasible paths generated by RC-S. In addition, the number of available robots $n_{r}$ also limits the upper bound of occupied VPs. Above all, the expression for calculating $n^{occupied}_{VP}$ can be formulated as:
$n^{occupied}_{VP}=\min\{\kappa\cdot\beta\cdot n_{VP},n_{r}\}$ (13)
@@
Now we investigate the forms of the two factors.
@@
This factor represents the impact of the number of workers on the quantity of VPs that can be occupied. For each aisle, the newly generated VPs at the entrance could be occupied by loaded robots when its corresponding loading station is active and worker is engaged in tasks. The quantity of such VPs is a critical determinant of the maximum throughput. Consider, for instance, a scenario where eight workers are stationed in an RSS with a $6+6$ aisle network. To ensure the balance of traffic flow and mitigate congestion at the aisle network’s periphery, workers are are evenly divided into four groups and positioned along the four edges of the site, as depicted in Figure 9. Four horizontal aisles and four vertical aisles are served by workers. We merge the area covered by these aisles and obtain the blue polygon.
Figure 9: The impact of number of workers on the VPs that can be occupied
For ease of discussion, the notation $\alpha$ will be employed to denote the ratio of workers to the network’s maximum worker capacity.
@@
The proportion of the blue region, noted as the workforce factor $\kappa$, could be obtained by:
@@
If either (i) the candidate feasible paths in RC-S include no more than two turns, or (ii) the average travel distance of a task is less than $(2-\alpha)min\{n_{v},n_{h}\}$, then an upper bound of the proportion of occupied VPs is $\kappa$;
The proof is presented in Online Appendix B. Proposition 1 provides with support for estimating the maximized network throughput. Next, we analyze the utilization rate of VPs using the RC-S scheme.
@@
This factor represents the utilization of VPs by the feasible paths generated by RC-S. We begin with examine the path reservation strategy within the framework of RC-S . In a first-come-first-served basis, the system allocates to each newly requested sorting task an available robot, assigning it the currently shortest viable route. This robot then proceeds through the network along a predefined spatio-temporal path, ensuring a collision- and queue-free delivery within the aisle network. However, this spatio-temporal reservation mechanism obligates the robot to adhere strictly to its assigned spatio-temporal trajectory, precluding any alterations like deceleration or temporary halts. Such inflexibility can lead to underutilization of traffic capacity, as VPs that are momentarily unoccupied cannot always be sequentially linked to construct a continuous path if they do not intersect at the same point in time. To quantify the impact of this reduction, we introduce an attenuation factor $\beta$, representing the spatial utilization.
Figure 10: The ”pile up” process of spatiotemporal occupation under RC
Figure 11: The loss of capacity in different network
@@
The attenuation factor $\beta$ rises with the increase of $(n_{h}+n_{v})$. Under the RC-S protocol, the aisle’s spatiotemporal occupation showcases a ”pile-up” effect, as depicted in Figure 10. The figure’s horizontal axis marks the spatial position within an aisle, while the vertical axis tracks discretized time intervals. Each robot’s transit through this path necessitates occupying a VP, visualized as a blue bar in the illustration. The collective occupation by different robots generates the ”pile-up” phenomenon. The capacity loss at any moment is represented by the unoccupied area beneath the upper boundary of this accumulation.
@@
When comparing networks of varying scales, i.e., those with shorter versus longer aisle lengths, we maintain a continuous approach: assuming that new sorting tasks emerge with starting points evenly spaced along the boundary and target destinations uniformly spread across the network. In a stable operation, the flow distribution of paths under consistent control logic remains analogous across different network sizes, with variations only in total path lengths. Accordingly, the pattern of unoccupied spaces beneath the upper boundary remains consistent across scales. Figure 11 illustrates this by comparing capacity losses across aisles of diverse lengths. Longer aisles tend to exhibit extended periods of spatiotemporal occupation, potentially exacerbating capacity loss — the unoccupied intervals shown in Figure 11. In consequence, we can reasonably infer that the absolute amount of lost capacity increases quadratically with the number of aisles $(n_{h}+n_{v})$, and the proportion of the lost traffic capacity should grow linearly (because the number of available VPs increases linearly with the length of aisle). In summary, the expression of attenuation factor $\beta$ is as follows:
@@
where $a$ and $b$ are two parameters. Parameter $a$ is slightly greater than 1 and $b$ is relatively small, indicating that for a small-scale network, the attenuation factor should be close to 1 because there are fewer conflict points in one feasible path. We use simulation to obtain the real performance of RC-S and apply linear regression to calibrate the value of the parameters in the attenuation coefficient. Experiments show that a proper value is $a=1.4$, $b=0.012$. Note that, due to the presence of the continuity assumption, there is a certain bias in $\beta$ when the number of aisles is small. Specifically, in smaller networks, this factor may overestimate the system’s throughput, especially in scenarios where the activation rate of loading stations is low. This is because some path shapes may not exist in such networks. Experimental validation of this effect will be presented in the subsequent section. To this end, we derive the final expression of number of occupied VPs:
$n^{occupied}_{VP}(\alpha,n_{r})=\min\{\frac{1-{(1-\alpha)}^{2}}{a+b(n_{h}+n_{v})}\cdot n_{h}n_{v},\ n_{r}\}$ (18)
@@
$n^{occupied}_{VP}(\alpha,n_{r})$ is non-decreasing and concave for all $\alpha$ and $n_{r}$ in its domain.
The proof of Lemma 1 can be accomplished by verifying that the partial derivatives of the two terms in the minimum function are non-negative and the Hessian matrices are both semi-negative definite. Lemma 1 indicates that opening more loading stations will increase the capacity of VPs within the network, thereby raising the upper limit of total throughput. However, the marginal returns of constructing new loading stations will gradually diminish.
@@
We now derive the average travel distance $\bar{l}$ of a sorting task, as the remaining part of the throughput estimation formula. Based on the mechanism of RC-S, an appropriate selection of paths is based on the principles of evenly distributed sorting demand, workload balancing of workers, and minimizing travel distance. Workload balancing principle ensures that loading stations do not experience starvation, thus preserving the stability of the system. It also aligns better with real-world scenarios. To maintain a balanced flow of robots between loading stations, we arrange the workers at the centers of each side, covering a length equal to $\alpha$ times the side length. Figure 12 illustrates this layout, where outlets are divided into four types of areas based on their positions. Sorting demand are uniformly and continuously distributed across the entire area. As mentioned earlier, each time a robot turns, it occupies additional space in the VP fleet, so it is necessary to restrict the maximum permissible robot’s turning times in one delivery. However, from the diagram, it can be observed that to ensure the outlets in area 3 are reachable from each loading station, paths with at least three turns are required. According to our simulation experiments, introducing paths with four or more turns did not effectively improve the sorting throughput. Therefore, we prioritize paths with fewer turns and do not consider paths that require more than three turns. We denote the path length serving area $i$ as $l_{i}$, with the proportion of this area represented by $p_{i}$. We will next present the rationale for their calculation and the formulas. For a detailed derivation process, please refer to the appendix B.
Figure 12: Dividing areas according to the layout
l
@@
Firstly, for area 1, servicing the outlets requires one change of direction, as shown in Figure 13(a1). Based on the balance of workload, the turning points should have an equal probability of occurring within the central area. We then perform a weighted accumulation based on the length and then take the average. Consequently, we obtain expressions for the area proportion and the expected path length:
@@
$E[l_{1}]$ $=2D\cdot[\frac{\frac{9+a^{2}}{6}(n_{h}^{2}+n_{v}^{2})-n_{h}n_{v}-\frac{1}{3}}{n_{h}+n_{v}}+1]$ (20)
For paths that pass by the areas of area 2, their endpoints are located at either the lower or upper loading stations, as shown in Figure 13(b1) and (b2), respectively. Let $(x_{1},y_{1})$, $(x_{2},y_{2})$ denote the coordinates of the two turning points in a path. $x_{1}$ and $x_{2}$ are uniformly distributed over the length covered by the workers, while the $y_{1}$ and $y_{2}$ are identical and uniformly distributed across the area. Notably, when x1 = x2, the path actually does not involve any turns. Let the robot choose the nearest top or bottom exit point when making the second turn. The expressions for $p_{2}$ and $E[l_{2}]$ are as follows:
@@
$E[l_{2}]$ $=2D\cdot[\frac{\alpha(n_{h}^{2}+n_{v}^{2})}{3(n_{h}+n_{v})}-\frac{2}{3\alpha(n_{h}+n_{v})}+\frac{3n_{h}n_{v}}{2(n_{h}+n_{v})}]$ (22)
For Area 3, as previously mentioned, only paths involving three turns can service it. Area 3 includes four individual small areas. The two small areas on the bottom, as their service paths shown in Figure 13(c1), have a length equal to the paths with one turn in Figure 13(a1); the two areas on the top, with service paths depicted in Figure 13(c2), involve two extra segments of path in the y-direction (or in the x-direction when the starting point is on the left or right side). As a result, the expressions for $p_{3}$ and $E[l_{3}]$ are as follows:
@@
$E[l_{3}]$ $=2D\cdot(\frac{n_{h}+n_{v}}{2}+\frac{(1+\alpha)}{4}\cdot\frac{n_{h}n_{v}}{n_{h}+n_{v}})$ (24)
Finally, for Area 4, which is in the center, we assume it is equally served by the paths of area 1 and area 2.
@@
Now, we further investigate the properties of equation (27).

### v4

@@ 4.2 Concept of RC-S
The entry cycle of a specific feasible path is the earliest cycle in which all VPs in the path are unoccupied.
A feasible path corresponds to a specific spatio-temporal trajectory pattern, represented by the sequence of VP reservations it occupies. The entry cycle denotes the earliest possible release cycle for the trajectory pattern, given the current VP reservation status.
Before the beginning of each cycle, the RC-S scheme performs path planning for all loaded robots that are waiting at entrance nodes and do not yet have assigned paths. These robots execute their sorting tasks based on the spatio-temporal trajectories specified by their assigned feasible paths. The corresponding feasible path assignment (FPA) problem is formulated as an integer programming model with binary decision variables. The objective is to minimize the total completion time of all unassigned tasks in the current cycle, together with penalties for any task that cannot be assigned a feasible path.
To reduce computational complexity, VP reservations from previously assigned paths are treated as fixed input. Consequently, the current planning process is based on the real-time availability of VP resources, leading to a greedy insertion approach. The main notation used in the FPA model is listed in Table LABEL:Notataion_fpa.
Table 2: Notations and explanation
Notations Explanation
Sets
$\mathcal{V}$ set of all nodes
$\mathcal{L}$ set of all loading stations
$\mathcal{O}$ set of all outlets
$\mathcal{C}$ set of all cycles starting from the current cycle
$\mathcal{R}_{i,k}$ set of feasible path patterns starting from loading station $i\in\mathcal{L}$ and passing by outlet $k\in\mathcal{O}$, independent of starting cycle
$\mathcal{R}_{i}^{(0)}$ Set of feasible paths available to loading station $i\in\mathcal{L}$ at the beginning of current cycle, based on the current destination outlet of the pending sorting task. If no task is pending, $\mathcal{R}_{i}^{(0)}=\emptyset$
$\hat{\mathcal{R}}_{i}$ set of feasible paths ending at loading station $i\in\mathcal{L}$
Parameters
$c_{i}^{r}$ completion time of feasible path $r\in\mathcal{R}_{i}^{(0)}$, measured in number of cycles
$\hat{c}_{i}$ penalty for delaying the assignment of a feasible path to a sorting task from loading station $i\in\mathcal{L}$ until the next cycle
$\delta_{i}^{r,t,\nu,l}$ incidence between feasible path $r\in\mathcal{R}_{i}^{(0)}$ starting at cycle $t\in\mathcal{C}$, node $\nu\in\mathcal{V}$ and cycle $l\in\mathcal{C}$
$N_{\nu}^{l}$ remaining capacity in node $\nu\in\mathcal{V}$ in cycle $l\in\mathcal{C}$
$d_{i}$ sorting demand from loading station $i\in\mathcal{L}$, $d_{i}\in\{0,1\}$
Variables
$x_{i}^{r,t}$ decision on whether feasible path $r\in\mathcal{R}_{i}^{(0)}$ starting at cycle $t\in\mathcal{C}$ is reserved
$\hat{x}_{i}$ decision on whether the sorting task from loading station $i\in\mathcal{L}$ should be postponed to the next cycle
$\underset{x_{i}^{r,t},\hat{x}_{i}}{\min}\ \sum_{i\in\mathcal{L}}[\sum_{r\in\mathcal{R}_{i}^{(0)}}\sum_{t\in\mathcal{C}}(c_{i}^{r}+t)x_{i}^{r,t}+\hat{c}_{i}\hat{x}_{i}]$
$s.t.\quad$ $\sum_{i\in\mathcal{L}}\sum_{r\in\mathcal{R}_{i}^{(0)}}\sum_{t\in\mathcal{C}}\delta_{i}^{r,t,\nu,l}x_{i}^{r,t}\leq N_{\nu}^{l}$ $\forall\nu\in\mathcal{V},l\in\mathcal{C}$ (9)
$\hat{x}_{i}+\sum_{r\in\mathcal{R}_{i}^{(0)}}\sum_{t\in\mathcal{C}}x_{i}^{r,t}=d_{i}$ $\forall i\in\mathcal{L}$ (10)
$x_{i}^{r,t},\hat{x}_{i}\in\{0,1\}$ $\forall i\in\mathcal{L},r\in\mathcal{R}_{i}^{(0)},t\in\mathcal{C}$ (11)
We introduce two sets of binary decision variables. Variable $x_{i}^{r,t}$ equals one if the feasible path $r\in\mathcal{R}_{i}^{(0)}$ starting at cycle $t$ is assigned to the robot at loading station $i$, and zero otherwise. Variable $\hat{x}_{i}$ equals one if this robot is deferred to the next cycle, and zero otherwise. $\mathcal{R}_{i}^{(0)}$ denotes the set of all feasible paths available to loading station $i$ in the current cycle $t_{0}$. If a sorting task is pending with destination outlet $k$, then $\mathcal{R}_{i}^{(0)}=\mathcal{R}_{i,k}$; otherwise, $\mathcal{R}_{i}^{(0)}=\emptyset$. In the objective function, $c_{i}^{r}$ denotes the completion time of path $r$, and $\hat{c}_{i}$ denotes the penalty for delaying the assignment of a feasible path to a loaded robot until the next cycle. This penalty is introduced to prevent robots from being assigned paths with excessively long completion times. Constraints (9) enforce the spatio-temporal occupancy limits of each node $\nu$, where $N_{\nu}^{l}$ denotes the remaining capacity of node $\nu$ in cycle $l$. $N_{\nu}^{l}$ is dynamically updated based on the FPA decisions in previous cycles. Conflict nodes are allowed to accommodate up to two robots per cycle; all other nodes are restricted to one. The parameter $\delta_{i}^{r,t,\nu,l}$ equals one if feasible path $r$ starting at cycle $t$ will occupy node $\nu$ in cycle $l$, and zero otherwise. Constraints (10) ensure that exactly one assignment is made for each sorting task, where $d_{i}$ equals one if there is a sorting task pending at loading station $i$. Finally, Constraints (11) impose integrality conditions.
It is essential to solve the FPA problem efficiently within each cycle. In large-scale networks, the number of feasible paths grows rapidly, making it challenging to obtain a solution within a reasonable computation time. We propose a heuristic algorithm to generate a feasible solution in a relatively short time, as shown in Algorithm 1.
Algorithm 1 Heuristic Algorithm for Implementing RC-S
0:  Initialize the reservation table that records the status of VP reservation for the upcoming $N_{record}$ cycles, labeled as unoccupied. Set the waiting time of each robot that is waiting to be assigned a feasible path as 0. Set the maximum search range in each cycle as $N_{search}$. Pre-compute the set of feasible path patterns with less than three turns as $\mathcal{R}=\bigcup_{i\in\mathcal{L},\ k\in\mathcal{O}}\mathcal{R}_{i,k}$. Initialize the candidate robot list as $\mathcal{S}_{r}=\emptyset$. Set cycle index $l=1$.
@@
0:  At the beginning of cycle $l$, for each loading station, if the first robot in the queue has not been assigned a feasible path, add it to the candidate list $\mathcal{S}_{r}$.
0:  If $\mathcal{S}_{r}$ is empty, go to Step 5. Otherwise, select the robot $a$ with the longest waiting time from $\mathcal{S}_{r}$. Obtain its origin loading station $i$ and target outlet $k$. Search the shortest feasible path with the earliest entry cycle for robot $a$: for future cycle $t\in[l,l+N_{\text{search}}]$:    for feasible path $r\in\mathcal{R}_{i,k}$ (sorted by path duration in ascending order):       Check if reserving path $r$ starting at cycle $t$ is feasible under current reservation table.      If feasible, record $(r,t)$ as the assigned path for robot $a$, and go to Step 3.   end forend for
0:  If a feasible pair $(r,t)$ was recorded in Step 2, then the VP-cycle occupation of this assignment is given by $\{(p_{i},l_{i}+t)\}$. Update the reservation table with $\{(p_{i},l_{i}+t)\}$, and assign the path to robot $a$. If no such pair is found, increase the waiting time of robot $a$ by $\tau_{c}$.
0:  Remove robot $a$ from $\mathcal{S}_{r}$, and return to Step 2.
@@
0:  Add the newly generated VPs in the next cycle into the reservation table, labeled as unoccupied. Remove the VPs that have exited the network.
0:  Set $l=l+1$. Update the status of all robots and assign parcels to empty robots queueing in loading zone. Return to Step 1.
In Algorithm 1, we introduce a heuristic search method that iteratively identifies the shortest feasible path with the earliest entry cycle, as a substitute for solving the original FPA model. During initialization, the reservation table is constructed with dimensions $n_{h}\times n_{v}\times N_{record}$, since the aisle network consistently contains $n_{h}\times n_{v}$ VPs at any given time. The value of $N_{record}$ is required to be greater than the maximum duration of feasible paths to fully capture the impact of each assignment on future spatio-temporal resource availability.
For each feasible path, we retain only the earliest admissible release cycle, as defined in Definition 2. By introducing a maximum search range $N_{search}$ in Step 3, we ensure that the number of pairs of feasible paths and entry cycles is finite, with an upper bound equal to the number of trajectory patterns $\lvert\mathcal{R}\rvert$ multiplied by $N_{search}$.
In each cycle, VPs move according to the RC-S pattern, accompanied by the generation of new VPs and the departure of others. The reservation table must be updated accordingly, as shown in Step 5. In Step 6, the path-assignment results for the current cycle are executed by the robots, which then perform movement, parcel drop-off, queueing, and loading within the cycle. The algorithm then proceeds to the planning phase of the next cycle.
[…]
@@ 5 Performance Analysis of RC-S
Two important metrics in traffic analysis (DAGANZO2010434) are the total vehicular distance traveled per second of operation, $\bar{m}(veh\cdot m/s)$, and the average travel distance, $\bar{l}$. The relationship between system throughput $\tilde{T}_{O}$ and these two metrics is shown in Equation (12):
@@
Under the mechanism of RC-S, robots are required to follow VPs for movement, so their average speed within the sorting zone is approximately equal to $v_{VP}$. Although acceleration and deceleration inevitably occur during turning, we limit the maximum number of turns in each feasible path and mitigate their impact on the average robot speed through the design of the robot turning and VP transfer process, as described in Section 4. Therefore, the first metric $\bar{m}$ can be estimated as the product of the number of occupied VPs, denoted by $n^{\text{occupied}}_{VP}$, and the speed of each VP, denoted by $v_{VP}$:
$\bar{m}\approx n^{\text{occupied}}_{VP}\cdot v_{VP}$ (13)
The speed of each VP can be obtained from Expressions (6)–(7), given the parameters of RC-S. For the number of occupied VPs, an intuitive assumption is that it depends on the number of workers engaged in loading parcels in the loading zone and on the spatial utilization under RC-S. We therefore introduce two factors: the workforce factor $\kappa$, representing the impact of the number of workers on the quantity of VPs that can be occupied; and the attenuation factor $\beta$, representing the utilization of VPs by the feasible paths generated by RC-S. In addition, the number of available robots $n_{r}$ also limits the maximum number of occupied VPs. Combining these effects, the expression for calculating $n^{\text{occupied}}_{VP}$ can be formulated as:
$n^{\text{occupied}}_{VP}=\min\left\{\kappa\cdot\beta\cdot n_{VP},n_{r}\right\}$ (14)
@@
We now investigate the forms of the two factors $\kappa$ and $\beta$.
@@
The workforce factor $\kappa$ represents the impact of the number of workers on the number of VPs that can be occupied. It is determined by the network size $(n_{h},n_{v})$ and the number of workers $n_{w}$. For each aisle, the newly generated VPs at the entrance can be occupied by loaded robots when the corresponding loading station is active and the worker at that station is engaged in tasks. The number of such VPs is a critical determinant of the maximum throughput. Consider, for instance, a scenario where eight workers are stationed in an RSS with a $6+6$ aisle network. To ensure balanced traffic flow and mitigate congestion at the periphery of the aisle network, the workers are evenly divided into four groups and positioned along the four edges of the site, as shown in Figure 10. Four horizontal aisles and four vertical aisles are served by these workers. We merge the areas covered by these aisles and obtain the blue polygon.
Figure 10: The impact of the number of workers on the VPs that can be occupied
For ease of discussion, we use $\alpha$ to denote the ratio of the number of workers to the network’s maximum worker capacity.
@@
The proportion of the blue area, referred to as the workforce factor $\kappa$, is derived by:
@@
If either (i) the candidate feasible paths in RC-S include no more than two turns, or (ii) the average travel distance of a task is less than $(2-\alpha)\min\{n_{v},n_{h}\}$, then an upper bound on the proportion of occupied VPs is $\kappa$.
The proof is presented in Online Appendix B. Proposition 2 provides theoretical support for estimating the maximum network throughput. Next, we analyze the utilization rate of VPs under the RC-S scheme.
@@
The attenuation factor $\beta$ represents the utilization of VPs by the feasible paths generated by RC-S. We begin by examining the path reservation strategy within the framework of RC-S. On a first-come-first-served basis, the system allocates an available robot to each newly requested sorting task, assigning it the currently shortest viable route. This robot then proceeds through the network along a predefined spatio-temporal path, ensuring collision- and queue-free delivery within the aisle network. However, this spatio-temporal reservation mechanism obligates the robot to adhere strictly to its assigned spatio-temporal trajectory, precluding any alterations such as deceleration or temporary halts. Such inflexibility can lead to underutilization of traffic capacity, as VPs that are momentarily unoccupied cannot always be sequentially linked to construct a continuous path if they do not intersect at the same point in time. To quantify the impact of this reduction, we introduce the attenuation factor $\beta$, representing the spatial utilization.
Figure 11: The ”pile-up” process of spatio-temporal occupation under RC
Figure 12: The loss of capacity in different networks
@@
The attenuation factor $\beta$ increases with $n_{h}+n_{v}$. Under the RC-S protocol, the spatio-temporal occupancy of an aisle exhibits a ”pile-up” effect, as depicted in Figure 11. The figure’s horizontal axis marks the spatial position within an aisle, while the vertical axis tracks discretized time intervals. Each robot’s transit through this path necessitates occupying a VP, visualized as a blue bar in the illustration. The collective occupancy by different robots generates the ”pile-up” phenomenon. The capacity loss at any moment is represented by the unoccupied area beneath the upper boundary of this accumulation.
@@
When comparing networks of varying scales, i.e., those with shorter versus longer aisle lengths, we adopt a continuous approach: we assume that new sorting tasks emerge with starting points evenly spaced along the boundary and target destinations uniformly spread across the network. In stable operation, the flow distribution of paths under consistent control logic remains analogous across different network sizes, with variations only in the total path lengths. Accordingly, the pattern of unoccupied spaces beneath the upper boundary remains consistent across scales. Figure 12 illustrates this by comparing capacity losses across aisles of different lengths. Longer aisles tend to exhibit extended periods of spatio-temporal occupancy, potentially exacerbating capacity loss–the unoccupied intervals shown in Figure 12. Consequently, we can reasonably infer that the absolute amount of lost capacity increases quadratically with the number of aisles $(n_{h}+n_{v})$, and that the proportion of the lost traffic capacity grows linearly (because the number of available VPs increases linearly with the length of the aisles). In summary, the expression of the attenuation factor $\beta$ is as follows:
@@
where $a$ and $b$ are two parameters. Parameter $a$ is slightly greater than 1 and $b$ is relatively small, indicating that for a small-scale network, the attenuation factor should be close to 1 because there are fewer conflict points in a feasible path. We use simulations to obtain the actual performance of RC-S and apply linear regression to calibrate the values of the parameters in the attenuation factor. Experiments show that appropriate values are $a=1.4$ and $b=0.012$. Note that, due to the continuity assumption, there is a certain bias in $\beta$ when the number of aisles is small. Specifically, in smaller networks, this factor may overestimate the system throughput, especially in scenarios where the activation rate of loading stations is low. This is because some path shapes may not exist in such networks. Experimental validation of this effect will be presented in the subsequent section. To this end, we derive the final expression for the number of occupied VPs:
$n^{\text{occupied}}_{VP}(\alpha,n_{r})=\min\left\{\frac{1-{(1-\alpha)}^{2}}{a+b(n_{h}+n_{v})}\cdot n_{h}n_{v},\ n_{r}\right\}$ (19)
@@
$n^{\text{occupied}}_{VP}(\alpha,n_{r})$ is non-decreasing and concave in both $\alpha$ and $n_{r}$ over its domain.
The proof of Lemma 1 can be established by verifying that the partial derivatives of the two terms in the minimum function are non-negative and that the Hessian matrices are both negative semi-definite. Lemma 1 indicates that opening more loading stations will increase the capacity of VPs within the network, thereby raising the upper limit of the total throughput. However, the marginal returns of constructing new loading stations will gradually diminish.
@@
We now derive the average travel distance $\bar{l}$ of a sorting task as the remaining component of the throughput estimation formula. Under the mechanism of RC-S, appropriate path selection follows the principles of evenly distributed sorting demand, balanced worker workload, and minimized travel distance. The workload-balancing principle ensures that loading stations do not experience starvation, thus preserving the stability of the system. It also aligns well with real-world scenarios. To maintain a balanced flow of robots between loading stations, we arrange the workers at the centers of each side, covering a length equal to $\alpha$ times the side length. Figure 13 illustrates this layout, where the sorting zone is divided into four types of areas based on the spatial locations of the outlets. Sorting demand is assumed to be uniformly and continuously distributed across the entire area. As mentioned earlier, each time a robot turns, it occupies additional space in the VP fleet, so it is necessary to restrict the maximum permissible number of turns in one delivery. However, from the diagram, it can be observed that to ensure the outlets in Area 3 are reachable from each loading station, paths with at least three turns are required. According to our simulation experiments, introducing paths with four or more turns does not effectively improve the sorting throughput. Therefore, we prioritize paths with fewer turns and do not consider paths that require more than three turns. We denote the path length serving Area $i$ by $l_{i}$, with the proportion of this area represented by $p_{i}$. We next present the rationale for their calculation and the corresponding formulas. For a detailed derivation, please refer to Appendix B.
Figure 13: Division of areas according to the layout
@@
First, for Area 1, servicing the outlets requires one change of direction, as shown in Figure 14(a1). Assuming balanced workload, the turning points should have equal probability of occurring within the central area. We then perform a weighted accumulation based on the length and take the average. Consequently, we obtain the following expressions for the area proportion and the expected path length:
@@
$E[l_{1}]$ $=2D\cdot\left[\frac{\frac{9+\alpha^{2}}{6}(n_{h}^{2}+n_{v}^{2})-n_{h}n_{v}-\frac{1}{3}}{n_{h}+n_{v}}+1\right]$ (21)
For paths that serve Area 2, their endpoints are located at either the lower or upper loading stations, as shown in Figure 14(b1) and (b2), respectively. Let $(x_{1},y_{1})$ and $(x_{2},y_{2})$ denote the coordinates of the two turning points in a path. $x_{1}$ and $x_{2}$ are uniformly distributed over the length covered by the workers, while $y_{1}$ and $y_{2}$ are identical and uniformly distributed across the area. Notably, when $x_{1}=x_{2}$, the path actually does not involve any turns. We assume that the robot chooses the nearest top or bottom exit point when making the second turn. The expressions for $p_{2}$ and $E[l_{2}]$ are as follows:
@@
$E[l_{2}]$ $=2D\cdot\left[\frac{\alpha(n_{h}^{2}+n_{v}^{2})}{3(n_{h}+n_{v})}-\frac{2}{3\alpha(n_{h}+n_{v})}+\frac{3n_{h}n_{v}}{2(n_{h}+n_{v})}\right]$ (23)
For Area 3, as previously mentioned, only paths involving three turns can serve it. Area 3 includes four individual small areas. The two small areas at the bottom, whose service paths are shown in Figure 14(c1), have a length equal to that of the paths with one turn in Figure 14(a1); the two areas at the top, with service paths depicted in Figure 14(c2), involve two extra segments of path in the $y$-direction (or in the $x$-direction when the starting point is on the left or right side). As a result, the expressions for $p_{3}$ and $E[l_{3}]$ are as follows:
@@
$E[l_{3}]$ $=2D\cdot\left(\frac{n_{h}+n_{v}}{2}+\frac{(1+\alpha)}{4}\cdot\frac{n_{h}n_{v}}{n_{h}+n_{v}}\right)$ (25)
Finally, for Area 4, which is in the center, we assume that it is equally served by the paths of Areas 1 and 2.
@@
We now further investigate the properties of Equation (28).